Circular No. 6338

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 B. Giles and T. Strohmayer, Goddard Space Flight Center and Univer-

sities Space Research Association, report: “The PCA experiment onboard XTE has now observed the galactic-center transient GRO J1744–28 (IAUC 6291) fourteen times for 1698 min during the period Jan. 18–Mar. 5. The nonbursting flux below 15 keV shows a linear decline since late January, which if continued will reach zero ∼ 1 week into April. This is a slower decrease than reported by BATSE for 20–100 keV (IAUC 6335). The flux in the main bursts is also decreasing, but at a slower rate, and will if continued reach zero ∼ 40–50 days later. A total of 50 large bursts has been seen (6to 8-fold increase over the persistent emission, 10–15 min exponential recovery). In addition, at least 5 ‘mini-bursts’ have occurred, where the flux increases by a factor of 2 and the recovery takes 20 s or less. The mini-burst duration appears slightly less than the more frequent, larger bursts. Many even smaller and shorter burst events often seem to be clustered in the few minutes prior to a large burst and are good predictors of the main events; these ‘micro-bursts’ last 1–2 s and show a 20- to 50-percent increase in countrate over the mean persistent rate. Finally, there are occasions when the flux appears to be primarily comprised of numerous spiky bursts of duration ∼ 1 s. This spiky variability is substantial compared to the statistical error over the 0.1- to 2-s timescale and to the ∼ 10-percent modulation at 2.2 Hz. The appearance of the similar bursting behavior (burst followed by a recovery in the flux) over such a wide range of amplitudes argues that a single phenomenon, such as accretion of plasma blobs of varying mass, may account for essentially all the source emission.”

α CENTAURI C E. Guinan and N. Morgan, Villanova University, report the determi-

nation of the rotation period of Proxima Cen (α Cen C) from the analysis of 3.5 months of monitoring with the International Ultraviolet Explorer during 1995 May-August: “Periodic variations in the strength of the Mg II h+k 280-nm line emission were found having a period of 31.5 ± 1.5 days and a mean amplitude of 20–25 percent. The observed modulation in the chromospheric Mg II emission arises as large active regions covering the dM4 star’s surface rotate in and out of view. Several flares were also observed. This rotation period is slightly faster than the 36.8 days found for α Cen B using the same method (see IAUC 6259).”

1996 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green